Solve the differential equation by separation of variables with step-by-step procedures.

(x+2)dx = (x+3) siny cosydy


1
Expert's answer
2021-09-17T00:06:40-0400
"(x+2)dx = (x+3) \sin y \cos ydy"

"\sin y \cos ydy=\dfrac{x+2}{x+3}dx"

Integrate


"\int \sin y \cos ydy=\int\dfrac{x+2}{x+3}dx"

"\int \sin y \cos ydy=\dfrac{1}{2}\int \sin (2y)dy"

"=-\dfrac{1}{4}\cos(2y)+C_1"



"\int\dfrac{x+2}{x+3}dx=\int\dfrac{x+3}{x+3}dx-\int\dfrac{1}{x+3}dx"

"=\int dx-\int\dfrac{1}{x+3}=x-\ln(|x+3|)+C_2"


"-\dfrac{1}{4}\cos(2y)+C_1=x-\ln(|x+3|)+C_2"

The solution of the equation in implicit form is


"x-\ln(|x+3|)+\dfrac{1}{4}\cos(2y)=C"